OpenStudy (anonymous):

Find the area of a sector with an arc length of 60 in. and a radius of 15 in. A. 117.75 in.² B. 282.60 in.² C. 450 in.² D. 6,750 in.² HELPPPPP!!!!!

OpenStudy (scarlettfarra2000):

l = theta/360 * 2Pi * r 60 = theta/360 * 2Pi * 15 theta/360 = 2/Pi theta = 720/Pi Area = theta/360 * Pi r^2 Area = 720/Pi * 1/360 * Pi 15^2 Area = 2*225 = 450 ins^2 When angle is in radians, you can say any of the following: {key: L = arc Length, ϴ = central angle, r = radius, A = sector area} L = rϴ A = ½r²ϴ or, combining the two above equations, we get a third gem: A = ½rL <--- this is the one we're going to use A = ½rL A = ½*60*15 = 450 in² [C]

OpenStudy (michele_laino):

first step, we have to find the subsequent angle \(\alpha\):